# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ Appendix E First-Order Semantics of the ANF Lazy $\lambda$-Calculus
The grammar includes both application let-bindings and value let-bindings (the latter bind a name to a value without application). Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results; ANF is neither a restriction nor an extension of the $\lambda$-calculus. Variables use de Bruijn indices: $\uparrow^{\!n}$ counts the number of enclosing scope levels (record braces) between the reference and its binder, following the same convention as inheritance-calculus. In particular, a $\lambda$-bound variable that crosses $k$ intervening $\mathbf{let}$-binding scopes and $n$ intervening $\lambda$-abstraction scopes is written $\uparrow^{\!n+k}$, not $\uparrow^{\!n}$. Let-bound variables are referenced by their binding name $x$.
Evaluation tree
An evaluation tree has a single abstraction: every node can serve as function, argument, binding, and result at the same time. A $\mathtt{Path}$ is a sequence of labels $(\ell_{1},\ldots,\ell_{k})$ identifying a position in the mixin tree. We write $p$ for a path, $()$ for the root, $p\mathbin{\triangleright}\ell$ for extension, $\mathtt{init}(p)$ for the parent path, and $\mathtt{last}(p)$ for the final label.
AST
An ANF term is parsed into an AST that provides two functions at each path $p$:
  • $\mathtt{children}(p)$: the set of labels locally defined at $p$.
  • $\mathtt{refs}(p)$: the set of reference pairs $(n,\;\ell_{*})$, where $n$ is a de Bruijn index and $\ell_{*}$ is a list of downward projections.
Each ANF construct populates these as follows.
An abstraction $\lambda x.\,M$ at path $p$:
$\mathtt{children}(p)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$; the subtree at $p\mathbin{\triangleright}\mathrm{argument}$ is empty ($\mathtt{children}=\varnothing$, $\mathtt{refs}=\varnothing$).
A let-binding $\mathbf{let}\;x=V_{1}\;V_{2}\;\mathbf{in}\;M$ at path $p$:
$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ represents the application $V_{1}\;V_{2}$ and is determined by the form of $V_{1}$:
  • If $V_{1}$ is a variable (either $\uparrow^{\!m}$ or a let-bound name), then
    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\{(n_{1},\;\ell_{*})\},$
    where the reference $(n_{1},\;\ell_{*})$ encodes the translation of $V_{1}$ (see below).
  • If $V_{1}=\lambda y.\,M^{\prime}$, the $\lambda$-literal is inlined:
    $\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{argument},\;\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing,$
    and the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $M^{\prime}$.
In both cases the subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{argument}$ is determined by $V_{2}$.
A value let-binding $\mathbf{let}\;x=V\;\mathbf{in}\;M$ at path $p$:
$\mathtt{children}(p)$ $=\{x,\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is determined by $M$. The subtree at $p\mathbin{\triangleright}x$ wraps the value in a $\mathrm{result}$ scope:
$\mathtt{children}(p\mathbin{\triangleright}x)$ $=\{\mathrm{result}\},$ $\mathtt{refs}(p\mathbin{\triangleright}x)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}x\mathbin{\triangleright}\mathrm{result}$ is determined by $V$. The accessor $x.\mathrm{result}$ thus yields the translated value, matching the convention for application let-bindings.
A tail call $V_{1}\;V_{2}$ at path $p$:
$\mathtt{children}(p)$ $=\{\mathrm{tailCall},\;\mathrm{result}\},$ $\mathtt{refs}(p)$ $=\varnothing.$
The subtree at $p\mathbin{\triangleright}\mathrm{tailCall}$ is analogous to the application node $p\mathbin{\triangleright}x$ above: if $V_{1}$ is a variable, $\mathtt{children}=\{\mathrm{argument}\}$ with $\mathtt{refs}=\{(n_{1},\;\ell_{*})\}$; if $V_{1}$ is a $\lambda$-literal, $\mathtt{children}=\{\mathrm{argument},\;\mathrm{result}\}$ with $\mathtt{refs}=\varnothing$ and the $\lambda$-body inlined. The subtree at $p\mathbin{\triangleright}\mathrm{result}$ is a forwarding reference:
$\mathtt{children}(p\mathbin{\triangleright}\mathrm{result})$ $=\varnothing,$ $\mathtt{refs}(p\mathbin{\triangleright}\mathrm{result})$ $=\{(0,\;(\mathrm{tailCall},\;\mathrm{result}))\}.$
Here $(0,\;(\mathrm{tailCall},\;\mathrm{result}))$ navigates to the enclosing scope ($p$ itself, since $n=0$ from $p\mathbin{\triangleright}\mathrm{result}$’s perspective), then projects $\mathrm{tailCall}.\mathrm{result}$.
A $\lambda$-bound variable $\uparrow^{\!n}$ used as an inheritance source at an application node $p$ contributes $(n,\;(\mathrm{argument}))$ to $\mathtt{refs}(p)$: navigate up $n$ scope levels, then project $\mathrm{argument}$. The same applies when $\uparrow^{\!n}$ occurs as a bare value in tail position or as an argument value, contributing the reference to $\mathtt{refs}$ at the corresponding path.
A let-bound variable $x$ at path $p$ is resolved lexically: find the nearest ancestor $p^{\prime}$ of $p$ such that $x\in\mathtt{children}(p^{\prime})$, let $n=|p|-|p^{\prime}|-1$, and contribute $(n,\;(x,\;\mathrm{result}))$ to $\mathtt{refs}(p)$, projecting $x.\mathrm{result}$ from the enclosing scope.
Note that since the de Bruijn index $n$ counts all scope levels (not just $\lambda$-abstractions), it accounts for intervening $\mathbf{let}$-binding and tail-call scopes. For example, in $\lambda.\,\lambda.\,\mathbf{let}\;r=\uparrow^{\!2}\;\uparrow^{\!0}\;\mathbf{in}\;r$, the outer-$\lambda$ variable has de Bruijn index $2$ (crossing the $\mathbf{let}$-scope and the inner-$\lambda$ scope), not $1$.
Both $\mathtt{children}$ and $\mathtt{refs}$ are pure data extracted from the AST; they are not functions of runtime state.
Labels
The labels observable at a path are those locally defined at any graft of any transitive callee:
(7) $\mathtt{labels}(p)=\bigl\{\;\ell\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}}),\;\ell\in\mathtt{children}(p_{\mathrm{graft}})\;\bigr\}$
Grafts
When a function is applied, its subtree is grafted at the call site. A child label $\ell$ of the callee also becomes a child of the caller; the child at the callee’s $\ell$ position and the child at the caller’s $\ell$ position (if any) share the same identity. The grafts of $p$ collect all such identity-sharing positions:
(8) $\mathtt{grafts}(p)=\begin{cases}\{p\}&\text{if }p=()\\[6.0pt] \{p\}\;\cup\;\left\{\;p_{\mathrm{callee}}\mathbin{\triangleright}\mathtt{last}(p)\;\left|\;\begin{aligned} &p_{\mathrm{callee}}\in\mathtt{callee}^{*}(\mathtt{init}(p)),\\ &\text{s.t.}\;\mathtt{last}(p)\in\mathtt{children}(p_{\mathrm{callee}})\end{aligned}\right.\right\}&\text{if }p\neq()\end{cases}$
Here $\mathtt{callee}^{*}$ denotes the reflexive-transitive closure of $\mathtt{callee}$, defined next.
Callee
The callee of a path $p$ is the set of paths that $p$ is grafted from. Concretely, $\mathtt{callee}$ resolves every reference in $p$’s grafts one step:
[…]
where $\ell_{*}=(\ell_{1},\ldots,\ell_{k})$. The $\mathtt{init}(p_{\mathrm{graft}})$ passes the reference’s enclosing scope to $\mathtt{scope}$, matching the convention of reference resolution (5), which calls $\mathtt{this}(\{p_{\mathrm{site}}\},\;\mathtt{init}(p_{\mathrm{def}}),\;n)$.
Scope resolution
$\mathtt{scope}$ navigates upward through the evaluation tree to find the binding site of a de Bruijn reference. Each ANF construct ($\lambda$-abstraction, $\mathbf{let}$-binding, or tail call) introduces one scope level; crossing a scope boundary means finding the call site that grafted the current subtree.
$\mathtt{scope}$ takes three arguments: an evaluation-site path $p_{\mathrm{site}}$ (where the grafted subtree lives), a definition-site path $p_{\mathrm{def}}$ (the position in the original AST), and a de Bruijn index $n$.
(10) $\mathtt{scope}(p_{\mathrm{site}},\;p_{\mathrm{def}},\;n)=\begin{cases}p_{\mathrm{site}}&\text{if }n=0\\[6.0pt] \mathtt{scope}\!\left(p_{\mathrm{caller}},\;\mathtt{init}(p_{\mathrm{def}}),\;n-1\right)&\text{if }n>0\end{cases}$
where $p_{\mathrm{caller}}$ is the unique path satisfying $(p_{\mathrm{caller}},\;p_{\mathrm{def}})\in\mathtt{callee}^{\mathrm{ctx}}(p_{\mathrm{site}})$, and $\mathtt{callee}^{\mathrm{ctx}}$ pairs each element of $\mathtt{callee}^{*}(p_{\mathrm{site}})$ with its provenance:
(11) $\mathtt{callee}^{\mathrm{ctx}}(p)=\bigl\{\;(\mathtt{init}(p_{\mathrm{step}}),\;p_{\mathrm{graft}})\;\big|\;p_{\mathrm{step}}\in\mathtt{callee}^{*}(p),\;p_{\mathrm{graft}}\in\mathtt{grafts}(p_{\mathrm{step}})\;\bigr\}$
At each step, $\mathtt{init}$ shortens $p_{\mathrm{def}}$ by one label and $n$ decreases by one. Since $n$ is a nonnegative integer, the recursion terminates.
This completes the chain of definitions needed to compute $\mathtt{labels}(p)$.
Observational fixpoint semantics
The definitions above constitute the complete semantics. Computation arises from querying progressively deeper paths in a lazily constructed tree: the observer drives computation by choosing which path to inspect, and the mutually recursive definitions unfold on demand.
Formally, the mutually recursive equations (7)–(11) define set membership recursively: $\ell\in\mathtt{labels}(p)$ holds if and only if it can be established by a finite chain of applications of these equations. Well-definedness follows from the same termination argument as for the mixin tree equations (Appendix A): if the recursive evaluation terminates, the reachable dependency subgraph is finite and acyclic, and the equations have a unique solution. When no finite derivation exists for a given query, the recursive evaluation does not terminate, analogous to divergence.
Convergence
An ANF term $M$ converges, written $M{\Downarrow}$, if there exists $n\geq 0$ such that
$\{\mathrm{argument},\,\mathrm{result}\}\;\subseteq\;\mathtt{labels}\!\bigl(\,\underbrace{\mathrm{result}\mathbin{\triangleright}\cdots\mathbin{\triangleright}\mathrm{result}}_{n}\,\bigr)$
in the semantics defined by the equations above. This corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: $M$ converges iff it has a head normal form (Theorem 4.2), and two terms are equivalent under this semantics iff they have the same Böhm tree (Theorem 4.3). The proofs, given in Appendix B, apply verbatim: each equation above is the specialization of the corresponding mixin tree equation (Section 3) to terms in the image of the translation $\mathcal{T}$ (Section 4), where the Single-Path Lemma (Lemma B.1) ensures that the set-valued $\mathtt{this}$ function (equation 6) degenerates to the single-valued $\mathtt{scope}$ (equation 10).
@@ F.2. Infrastructure
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$. This provides the right-hand (visitor) side of a join: when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
@@
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$. This provides the left-hand (visitee) side: the relation being iterated. Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
@@ F.8. Recursive Rules
Recursive Datalog rules, such as transitive closure $\mathrm{path}(X,Y)\mathrel{{:}\!{-}}\mathrm{edge}(X,Z),\;\mathrm{path}(Z,Y)$, are encoded identically to non-recursive rules: the head relation $\mathrm{path}$ appears both as a defined scope and as a reference in the body (providing $\mathrm{WithVisitorTail}$ for the $Z$-join). This should not be read as ordinary self-reference of a single-valued lazy program. Operationally, tabled evaluation (Section 3, Appendix A) computes $\mathrm{lfp}(T_{P})$: each query into $\mathrm{path}$ triggers further evaluation through the recursive body, and the trie union ($\mathtt{supers}$) accumulates all reachable tuples. This corresponds exactly to Naive evaluation (Bancilhon and Ramakrishnan, 1986) of the $T_{P}$ operator (van Emden and Kowalski, 1976): each iteration applies every rule to the current fact set and adds newly derived facts. Because Datalog-encoded relations have finitely many constants per query, the per-query answer domain is finite, and convergence is guaranteed by the ascending chain condition on a finite lattice. Semi-Naive evaluation (Bancilhon and Ramakrishnan, 1986), which restricts each iteration to rules that use at least one newly derived fact, is a natural optimisation left for future work.
@@ F.9. Semantic Correspondence with Standard Datalog
For example, given $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$ (a 3-cycle), the two-hop self-join $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ produces exactly $3$ pairs: $\{(a,c),\;(b,a),\;(c,b)\}$, matching standard Datalog. This is the behavior exhibited by RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026). When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple (a,c)(a,c)—not all of $\{a,b,c\}$.

### 16 Mar 2026 v7

@@ Appendix E First-Order Semantics of the ANF Fixpoint $\lambda$-Calculus
The grammar includes application let-bindings ($\mathbf{let}\;x=V_{1}\;V_{2}$), value let-bindings ($\mathbf{let}\;x=V$), tail calls ($V_{1}\;V_{2}$), and tail values ($V$). Variables use de Bruijn indices: $\uparrow^{\!n}$ counts only $\lambda$-abstractions between the reference and its binder, following the standard convention. Let-bound variables are referenced by their binding name $x$.
Two node types
Each position in the Böhm tree is either a scope node or an application node. These have distinct path types $\mathcal{S}$ and $\mathcal{A}$ and distinct first-order functions.
A scope node (path type $\mathcal{S}$) represents a body: a chain of let-bindings ending in a tail expression. A scope may or may not be a $\lambda$-abstraction (recorded by $\mathrm{is\_abstraction}$). Its edges are the names of application let-bindings $\{x_{1},\ldots,x_{k}\}$; value let-bindings do not create edges (see AST below). Navigation from a scope to a binding: $s\mathbin{\triangleright}x_{i}$ gives an application path $a\in\mathcal{A}$.
An application node (path type $\mathcal{A}$) represents a function application $V_{1}\;V_{2}$ bound by a $\mathbf{let}$. In the AST, $\mathtt{head\_ref}$ and $\mathtt{arg\_ref}$ are references (de Bruijn index plus projection, or an inline $\lambda$-literal); in the evaluated Böhm tree, $\mathtt{head}$ and $\mathtt{argument}$ lazily resolve these references to scope nodes. Navigation from an application to a scope: $\mathtt{head}(a)\in\mathcal{S}$ and $\mathtt{argument}(a)\in\mathcal{S}$.
The path types alternate strictly:
$\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,x_{i}\;}\mathcal{A}\xrightarrow{\;\mathtt{head}/\mathtt{argument}\;}\mathcal{S}\xrightarrow{\;\mathbin{\triangleright}\,y_{j}\;}\mathcal{A}\;\cdots$
There is no scope-to-scope or application-to-application direct transition.
@@ F.2. Infrastructure
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisitorTail}$. This provides the right-hand side of a join [^36] In the context of the visitor pattern, this acts as the visitor. : when two relations’ $\mathrm{VisitorMap}$ entries are composed, $\mathrm{VisitorTail}$ carries the joined relation’s row data for the matched constant.
@@
Maps each constant $d$ to its sub-trie entry $\mathrm{TailMap}.d$, labeled as $\mathrm{VisiteeTail}$. This provides the left-hand side [^37] This acts as the visitee. : the relation being iterated. Together, $\mathrm{WithVisitorTail}$ and $\mathrm{WithVisiteeTail}$ enable value-level joins where only constants present in both relations produce output.
@@ F.8. Recursive Rules
Recursive Datalog rules, such as transitive closure $\mathrm{path}(X,Y)\mathrel{{:}\!{-}}\mathrm{edge}(X,Z),\;\mathrm{path}(Z,Y)$, are encoded identically to non-recursive rules: the head relation $\mathrm{path}$ appears both as a defined scope and as a reference in the body (providing $\mathrm{WithVisitorTail}$ for the $Z$-join). This should not be read as ordinary $\mathtt{this}$ of a single-valued lazy program. Operationally, tabled evaluation (Section 3, Appendix A) computes $\mathrm{lfp}(T_{P})$: each query into $\mathrm{path}$ triggers further evaluation through the recursive body, and the trie union ($\mathtt{supers}$) accumulates all reachable tuples. This corresponds exactly to Naive evaluation (Bancilhon and Ramakrishnan, 1986) of the $T_{P}$ operator (van Emden and Kowalski, 1976): each iteration applies every rule to the current fact set and adds newly derived facts. Because Datalog-encoded relations have finitely many constants per query, the per-query answer domain is finite, and convergence is guaranteed by the ascending chain condition on a finite lattice. Semi-Naive evaluation (Bancilhon and Ramakrishnan, 1986), which restricts each iteration to rules that use at least one newly derived fact, is a natural optimisation left for future work.
The recursive rule as monadic pseudocode
The recursive rule above roughly corresponds to the following pseudocode (Haskell-like syntax with RebindableSyntax and OverloadedLists semantics, where do-notation desugars to a Set-based bind and list literals denote sets):
{-# LANGUAGE RebindableSyntax, OverloadedLists #-}
s >>= f = unions [f x | x <- s]
return x = [x]
guard True = [()]
guard False = []
edge = [("a","b"), ("b","c"), ("c","a")]
path = edge <> do
(x, y0) <- edge
(y1, z) <- path
guard (y0 == y1)
return (x, z)-}
Each line of the pseudocode has a direct counterpart in the inheritance-calculus encoding (Section F.4):
edge:
corresponds to the base-fact trie;
path = edge <> do …:
corresponds to $\mathrm{path}$ inheriting $\mathrm{edge}$ (base facts) and the recursive body (derived facts via $\mathtt{supers}$);
(x, y0) <- edge:
corresponds to Level 1 ($\mathrm{xAcc}$), iterating $X$ from $\mathrm{edge}$;
(y1, z) <- path:
corresponds to Level 2 ($\mathrm{yAcc}$), joining on $Y$ with $\mathrm{path}$, and Level 4 ($\mathrm{zAcc}$), iterating $Z$;
guard (y0 == y1):
corresponds to the $\mathrm{WithVisitorTail}\cap\mathrm{WithVisiteeTail}$ intersection in Level 2: a constant produces a $\mathrm{Visitor}$ entry only if it appears in both relations;
return (x, z):
corresponds to the $\mathrm{Replacement}$ chain in Level 5, constructing the output tuple $(X,Z)$.
The only difference is the evaluation strategy: Haskell evaluates path lazily (diverges), whereas inheritance-calculus evaluates it via tabled $T_{P}$ iteration (converges to $\mathrm{lfp}$).
However, this monadic program does not converge because path is defined recursively. [^38] Enabling the RecursiveDo extension does not help: mfix ties a knot at the element level via laziness, whereas the transitive closure equation requires a fixed point at the container level, that is, finding a set $S$ such that $S=F(S)$. Inheritance-calculus solves this through tabled evaluation (Section 3), which computes the least fixed point of the $T_{P}$ operator by iterating over the lattice of mixin trees, precisely the container-level fixed point that the pseudocode above cannot express.
@@ F.9. Semantic Correspondence with Standard Datalog
For example, given $\mathrm{edge}=\{(a,b),\;(b,c),\;(c,a)\}$ (a 3-cycle), the two-hop self-join $\mathrm{twoHop}(X,Z)\mathrel{{:}\!{-}}\mathrm{edge}(X,Y),\;\mathrm{edge}(Y,Z)$ produces exactly $3$ pairs: $\{(a,c),\;(b,a),\;(c,b)\}$, matching standard Datalog. This is the behavior exhibited by RelationalTwoHopRepeatedVisitor.mixin.yaml (Yang, 2026). When the $Y$-join matches constant $b$ (present in both $\mathrm{edge}$’s second column from $(a,b)$ and $\mathrm{edge}$’s first column), the $\mathrm{Visitor}$ for $b$ receives $\mathrm{VisitorTail}=\mathrm{edge}.\mathrm{TailMap}.b$ (containing only $c$), and $\mathrm{Replacement}$ constructs a single output tuple $(a,c)$, rather than all of $\{a,b,c\}$.